SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)
Activities, Problems, Nuts.....


See also action items-nonstandard connections here.

  • Operation Managers: Mario Pelliccioni, Antonio Petrella

  • Past Operation Managers:
    • July 2006: Alessandro Gaz, Mirco Andreotti
    • June 2006: Elisa Manoni, Alessandro Gaz
    • May 2006: Loredana Lopez, Elisa Manoni
    • April 2006: Loredana Lopez, Silvano Tosi
    • March 2006: Alessia D'Orazio, Silvano Tosi
    • January-February 2006: Emanuele Di Marco, Antonio Petrella
    • November-December 2005: Francesco Polci, Roberto Covarelli

    • July 23rd, 2006
    • Taken plateau measurement for Forward Endcap RPCs in streamer mode.
    • July 17th, 2006
    • Taken plateau measurement for Forward Endcap RPCs in avalanche mode.
    • June 26th, 2006
    • Cr4 CH #9_B and Cr5 CH #2_A tripped simultaneously quite a few times. It does not seem to be a problem in the power supply, so we move back to the previous configuration, switching the chambers on Cr5 CH #2_A to Cr4 CH #9_A.
    • June 22nd, 2006
    • Moved temperature alarm thresholds for PS of FEE_ECFE: yellow 56 -> 57, red 57 -> 59 degrees.
      Taken plateau measurement for Forward Endcap RPCs in avalanche mode.
    • June 20th, 2006
    • Taken plateau measurement for Forward Endcap RPCs in streamer mode.
    • June 6th, 2006
    • In order to check if the recent trips on Crate 4, channels #9_A, #9_B are due to a mis-behavior of PS4, we switched the chambers on C4 #9_A to C5 #2_A. The list of masked channels has been updated.
    • May 19th, 2006
    • Replaced the bottom FWD West Endcap bubbler because of an oil(704) leak.
    • May 17th, 2006
    • Avalanche RPCs plateau measurement.
    • May 16th, 2006
    • Forward Endcap plateau measurement: turned off both HV and LV for the avalanche RPCs during the measurement.
    • May 14th, 2006
    • Due to an outage of the cooling system of the racks in the electonic house, we lost all the off-line monitoring for the ifr (as happened for all the other subdetectors). When the problem of the cooling system was fixed, we noticed a strange feature in the DAQ system: even if the power outage was restored, the FEE creates were off. So we need to power cycle, from epics, the voltage for all the DAQ system.
    • May 8th, 2006
    • Tuning of the valve of bubbler #12 of Bottom Backward Endcap (it stopped bubbling on May 1st), but it's not bubbling anyway!
    • May 1st, 2006
    • Avalanche RPCs plateau measurement.
    • April 21th, 2006
    • Forward Endcap plateau measurement: turned off both HV and LV for the avalanche RPCs during the measurement.
    • April 10th, 2006
    • Access to the detector! Several things done:
      Replaced the power connector of minicrate B1 (see below).
      Took the voltage monitoring board of the avalanche preamps for repair.
      Re-connected the HV to FWD west top, layer 6.
      Switched avalanche chambers to the new gas mixture and took a plateau scan to find the new operating voltage.
    • April 9th, 2006
    • The power to the minicrate B1 for the FWD east RPCs died suddenly without giving any sign detectable by the shifters (and....not even an efficiency decrease!)
    • April 7th, 2006
    • The voltage read out in epics for the avalanche preamps started to show flacky values, oscillating between 0 and ~7V. We went inside and verified that the supplied voltage is the expected one (~5V): it's a readout problem.
    • March 31st, 2006
    • Replaced the pre-amplifiers of the top avalanche chambers.
    • March 29th, 2006
    • Forward Endcap plateau measurement.
      As we started to take cosmic runs,we saw that trigger rate was ~7000Hz and this would cause an high deadtime.
      We checked on single rate EPICS panel that this high rate was due to avalanche chambers.
      It's not sufficient to turn off the HV for this chambers to reduce the rate.
      The avalanche LV also needs to be turned off,because the high rate is due to noise electronic.
      When we finished to take plateau measurement, we left avalanche HV and LV off,
      to permit to take cosmic with all detectors without red alarm (caused by avalanche related rates).
    • March 28th, 2006
    • Since there was some works on the gas shack and the gas system has been turned off
      (see http://babar-hn.slac.stanford.edu:5090/HyperNews/get/DetOper/4679. html )
      the IFR has been rumped down (HV off) until the gas flow has been resumed.
    • March 27th, 2006
    • Replaced 5 FECs in the backward endcap (some were noisy, some were dead); 2 were replaced on the east side, 3 on the west side.
    • March 24rd, 2006
    • Turned back HV on during owl shift.
      IFR Mixed Gas Output Flow Fault Trip: SIAM 5 and SIAM 8 channels tripped. Resetted output fault and SIAMs trips on gas shack racks.
      Generally "output flow faults" are always spurious trips. The normal procedure is to simply re-set the trip.
      If it trips again within a short time then it would be better contact IFR GAS expert: it could be a real problem.
    • March 23rd, 2006
    • Turned back LV on after the test on the vesda system.
    • March 22nd, 2006
    • PEP shutdown starts: IFR put in safe mode for the door opening and SVT uncabling: HV turned off. Also LV turned off in light of a planned test on the vesda system.
      Adjusted flow on FEM5 to 30 cc/min. FEM5 flow was setted to 45 cc/min because we used it to take fluorine measurement.
      FEM5 gas output needs to be reconnected to bubbler.
    • March 21st, 2006
    • Selina Li (trigger expert) reports that the time alignment of the IFT object was shifted down lower as compared to other subdetectors' for about a week (run range 62822-63014). Investigation still ongoing: apparently in the same time period the avalanche chambers (layers 3 and 5 only) showed a 3-time larger than usual single rate.
    • March 19th, 2006
    • Another gas trip alarm: apparently a glitch. Reset by the shifter with no problem.
      Two channels started to show high occupancy in the fast-mon crate occupancy plot (channels 83 -masked- and 84 -unmasked-)
    • March 17th, 2006
    • Avalanche chamber bottle getting empty: switched to the other one.
    • March 12th, 2006
    • Gas trip alarm. Reset by the shifter with no problem.
    • March 09th, 2006
    • The gas bottle serving the avalanche RPCs has been replaced (there is also one more spare).
    • March 07th, 2006
    • We updated the reference histograms for the DQM.
      The gas bottle serving the avalanche RPCs is getting empty. Two new bottles have arrived and will be installed on March 9th. In the meanwhile the flux is set at 5 psi (central display reading); the online premix gas bottle pressure gauges reading is currently somewhat below 10 psi.
    • March 03rd, 2006
    • PS of the forward east that was problematic since a couple of days (it's the +7V one) replaced. The spare was already inside.
      After the replacement everything is fine! We also verified the possibility of controlling the new supply remotely
      (ability to turn it/off verified) and the test was successful.
      A new spare has been moved inside.
      Epics alarm level moved to default 5V.

      Looking at Ambient DB we see that PS voltage is 6.2V also if the Epics panel shows 6.8V.
      The problem is that IFR Archiver is dead. Problem has been fixed by experts.
      (see e-logbook here).

    • March 02nd, 2006
    • LV red alarms on forward East +7 V supply since yesterday(alarm paused run).
      Starting Mar 1 during the owl shift,Epics shows ~10 spikes down to 6V in the voltage.
      Today, within few hours( 10a.m.- 2p.m.)the voltage was consistently low:4-5V.
      Power cycling didn't help.
      We took an access to investigate. The supply voltage is 5.2V.We were unable to adjust upwards.
      It may be that the built in voltage limit is preventing us from further raising the voltage.
      The power supply must be replaced.
      Until the access the Epics alarm level has been lowered to 3V(from 5V) to allow PEP to inject.
      This,infact,will keep IFR flag runnable.Running with IFR 'ignored' was not OK because MCC could not trickle in.
      Firts run with lowered alarm threshold is 62645.
      The present voltage(~5V)may be enough to run the FECs as the scalar rates look nominal
      (but DQ should look at these runs closely....FE door is ~1/4 of the total muon solid angle).
    • flows on FEM 1,3,5 adjusted to 30,30,45 cc/respectively.

    • February 27th, 2006
    • we use FEM 5 and FEM 7 for the fluorine measurements(this means that on epics bubbler panel
      you see flow 0 because the gas output is connected to holes going into EH).
      Steve reconnected FEM2 and FEM3 to FEM2 and FEM3 bubbler channel,respectively.
      For the avalanche layers (FEM1,FEM3 and FEM5):
      FEM3 bubbler is going very quickly(~500) while FEM1 bubbler is ~100.
      the flow on FEM3 had been turned up so that it would bubble in the fluorine setup.
      Since it's no longer connected there we should turn its flow down so it is the same as FEM1.
      Instead, FEM5 should be turned up so that it's close to the top of rotameter gauge.
    • February 25th, 2006
    • HV in Crate 2, Channel #21 (accordingly to the maps should be RPC barrel in sxt 5 layer 19) started tripping
      continuosly during the closure of BaBar Forward doors.
      If we try to turn it on, the current jumps to 2000 immediately, causing a trip. It has been masked and turned off.
      It not seems to be a great loss since the efficiency for this layer has been 0 in the last month and also the current
      when the RPC is at V1 is very low (between 1mA and 3mA).
      We leave the chamber off and masked as long as we can have an access.
    • February 24th, 2006
    • The temperature of the lower tank chiller has been lowered from 13 degrees down to 12 degrees to stop
      the early-morning chiller trip we had for a couple of days. (probably caused by the bdg 620 door opened).
      We set the new temperature of alarm 1 on the regulator in IR2 (bldg 620) to 12 degrees.
      Then we also set the low alarm temperature of the chiller in EPICS accordingly to the regulator one.
    • January 13th, 2006
      • Forward Endcap plateau measurements
      • Avalanche chambers plateau measurements. Preliminary instructions here.
    • January 10th, 2006

    • During the access, we (Henry, Emanuele and Antonio)
    • Replugged the avalanche chambers layer 1 amplifiers, which were off since December due the oscillations.
    • Fixed a miscabling in minicrate E3
    • Fixed a loose connection for FEE on Forward EC, middle sector, bottom RPC.
    ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

    The same page for 2005 is here.